OpenStudy (anonymous):

this one should be some what fun Water flows into a tank according to the rate F(t) = (6 + t)/(1 + t), and at the same time empties out at the rate E(t) = (ln(t+2))/(t+1) , with both F(t) and E(t) measured in gallons per minute. How much water, to the nearest gallon, is in the tank at time t = 10 minutes. You must show your setup but can use your calculator for all evaluations.

ganeshie8 (ganeshie8):

presumably the tank is empty to start with..

OpenStudy (anonymous):

do i just evaluate both functions at t =10 and then subtract f(10) - E(10)? and Id assume the tank was empty i wasn't given anymore information

OpenStudy (irishboy123):

\(\frac{dV}{dt} = F(t) - E(t)\)

OpenStudy (irishboy123):

build and solve integral

OpenStudy (anonymous):

\[\int\limits \frac{ (t+6)-(\ln(t+2)) }{ (t+1) }\]?

ganeshie8 (ganeshie8):

Looks good, but that wont evaluate into elementary functions so you may simply use wolfram http://www.wolframalpha.com/input/?i=%5Cint_0%5E%2810%29+%5Cfrac%7B+%28t%2B6%29-%28%5Cln%28t%2B2%29%29+%7D%7B+%28t%2B1%29+%7D

OpenStudy (anonymous):

:( i was hoping it would be more complicated

ganeshie8 (ganeshie8):

evaluating it is complicated indeed but i think you should not waste time messing with special functions at this point..

OpenStudy (anonymous):

yeah your probably right , anyway thanks once again

ganeshie8 (ganeshie8):

np

ganeshie8 (ganeshie8):

maybe just double check the problem has no typoes because usually you should expect to see easily workable integrals in these type of problems.. the focus should be on setting up the integral, not evaluating it..

OpenStudy (anonymous):

okay one second,

OpenStudy (anonymous):

all seems clear to me , do you see anything out of place ?

ganeshie8 (ganeshie8):

integral becomes easy if E(t) is like below : \(E(t) = (\ln(t+\color{red}{1}))/(t+1)\)

OpenStudy (anonymous):

nope triple checked its t + 2

ganeshie8 (ganeshie8):

then we're good, nearest to gallon, the water in the tank is 18 gallon at t = 10

OpenStudy (anonymous):

@ganeshie8 can i ask you something?

ganeshie8 (ganeshie8):

sure ask

OpenStudy (anonymous):

why are you so helpfull? im serious you clearly are on a higher level of math, why do you keep helping people with silly linear calculus I problems, you couldn't possibly be learning anything and i'm sure your own work keeps you busy you have infinity medals what keeps you here ?

ganeshie8 (ganeshie8):

Haha thanks for the good words! But you have no idea, I suck at math, thats the only reason I'm still here ;p

OpenStudy (anonymous):

god , i must be truly awful then

ganeshie8 (ganeshie8):

I think you're doing great!

OpenStudy (anonymous):

:)